OpenStudy (anonymous):

Solve the given nonlinear inequality. Write the solution set using interval notation. x2 − 3x + 4/x + 1 ≤ 1

OpenStudy (anonymous):

\[x ^{2}3x+4/x+1\le1\]

OpenStudy (anonymous):

I guess that the inequality really is: \[x^2 - 3x + \frac {4}{x}+1\le1\] This is the same as: \[x^2-3x+\frac{4}{x}\le0\] I rewrite the equation \[x^2-3x+\frac{4}{x}=0\] by guessing the solution x=-1 and get the following: \[(x+1)(x^2-4)=0\] The ineq can be written as: \[(x+1)(x-2)(x+2)\le0\] So it is 0 at x=-2, x=-1 and x=+2 In the inequality will be less than zero when \[x \le-2\] or \[-1 \le x \le 2\] \[x \in [-\infty,-2]\cup[-1,2] \]